OpenStudy (anonymous):

what is -3=-6x x= and explain how and why thxs

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Your goal when given a variable is to isolate it. To do this, divide both sides by -6 resulting in 1/2 = x.

OpenStudy (anonymous):

-3=-6x 3/3=-6x/3 -1=-2x 1/2=-2x/2 -1/2=-x -(-1/2)=-(-x) 1/2=x

OpenStudy (anonymous):

1/2=-2x/2 is (-1/2=-2x/2)

OpenStudy (anonymous):

choezom your answer is incorrect